OpenStudy (anonymous):

There are 52 cards in a deck. There are 4 eights and 4 nines. What is the probability that if you cut the deck, you'll get an eight or a nine? (Points : 1)

OpenStudy (amorfide):

 i think... i would love for someone to correct me if im wrong

OpenStudy (anonymous):

2/13 or 8/52.Explanation, we think of the event of getting a 9card A=4/52, and 8card B=4/52, when we cut the deck, the events remain the same, A=B=2/26=4/52, so we have P(A U B)=8/52=2/13